# Beyond Diagonal RIS: Passive Maximum Ratio Transmission and Interference Nulling Enabler

### v1

@@ Abstract
Beyond diagonal reconfigurable intelligent surfaces (BD-RIS) generalizes and goes beyond conventional diagonal reconfigurable intelligent surfaces (D-RIS) by interconnecting elements to generate beyond diagonal scattering matrices, which significantly strengthen the wireless channels. In this work, we use BD-RIS for passive multiuser beamforming in multiuser multiple-input-single-output (MU-MISO) systems. Specifically, we design the scattering matrix of BD-RIS to either maximize the sum received signal power at the users following maximum ratio transmission (MRT), or to nullify the interference at the users following zero forcing (ZF). Furthermore, we investigate uniform/optimized power allocation and ZF precoding at the base station (BS). Numerical results show that BD-RIS improves the interference nulling capability and sum rate with fewer reflecting elements (REs) compared to D-RIS. In addition, at moderate to high signal to noise ratios (SNRs), passive interference nulling reduces the complexity at the BS by relaxing the need for precoding or water-filling power allocation design. Furthermore, the passive MRT with ZF precoding achieves a tight sum rate performance to the joint design considering MU-MISO scenarios with many REs while maintaining low computational complexity and simplifying the channel estimation.
@@ I-A Related Work
RIS technologies have gained significant attention recently due to their potential to enhance the communication performance of wireless systems. Among these technologies, two prominent categories have been intensively investigated in the literature, which are diagonal reconfigurable intelligent surfaces (D-RIS) [6, 7, 8, 9, 10, 11, 12] and beyond diagonal reconfigurable intelligent surfaces (BD-RIS) [13, 14, 15, 16, 17, 18, 19, 20, 21]. While these two categories are often discussed separately, it is important to note that BD-RIS can be viewed as a generalization of D-RIS. In other words, BD-RIS encompasses the characteristics of D-RIS as a special case, offering a broader and more flexible framework for reconfigurable intelligent surfaces. The key limitation in D-RIS is that it manipulates only the phase of the incident signal with $N$ impedances attached to $N$ REs making it single-connected. In contrast, BD-RIS manipulates both the phase and amplitude of the incident signal with a network of impedances that connect the $N$ REs together [13, 14]. By interconnecting all the REs to each other through $\frac{N(N+1)}{2}$ impedances, the fully-connected BD-RIS has been proposed, having the highest flexibility yet the highest circuit complexity. To balance BD-RIS circuit complexity and performance, group-connected architecture is introduced [13]. Also, graph theory is utilized to design optimal lower complexity architectures such as forest-connected and tree-connected [15]. While the above architectures are realized by reciprocal impedance networks, non-diagonal reconfigurable intelligent surfaces (ND-RIS) are proposed in [16] based on non-reciprocal impedance networks to increase the effective channel gain by ordering the channel gains in both directions.
@@
Focusing on various RIS architectures, in existing literature, beamforming design for both D-RIS [6, 7, 8, 9, 10, 11, 12] and BD-RIS [16, 17, 18, 19, 20, 21] aided multiuser multiple-input-single-output (MU-MISO) systems have been studied and explained in detail below.
@@
D-RIS: The authors of [6] minimize the transmit power at the base station (BS) while maintaining a certain signal to interference and noise ratio (SINR) level at the users by jointly optimizing the scattering matrix of D-RIS and the active precoder at the BS using the alternating optimization (AO) algorithm. Similarly, Yang et al. [7] achieve the same goal using passive maximum ratio transmission (MRT)/zero forcing (ZF) at the D-RIS while iteratively optimizing the power allocation at the BS. Considering the same setup, the authors of [8] use AO framework to maximize the minimum rate for a better user fairness. The authors of [9] derive a closed-form steering vectors-based multiuser linear precoder to null the known interference and minimize the unknown interference. Jiang et al. [10] use the AO framework to design the scattering matrix for interference nulling, achieving maximum degrees-of-freedom with high probability when the number of REs is $N>2K(K-1)$, with $K$ the number of users. The authors of [11] present a robust gradient-based meta-learning approach without pre-training to design the scattering matrix and the BS precoding matrix. Jangsher et al. [12] maximize the sum rate of a wireless unmanned aerial vehicle (UAV) with RIS considering energy efficiency constraints and practical limitations such as phase compensation and imperfect channel state information (CSI).
@@
BD-RIS: A two-stage design for the BD-RIS scattering matrix and BS precoding matrix is considered in [16, 17], unlike [18] which considers the joint design. For instance, Li et al. [16] consider ND-RIS with single-connected architecture and design the ND-RIS scattering matrix to maximize sum of the users’ combined channel gain following semi-define relaxation method, while the popular water-filling method is used for power allocation although interference has not been nullified. Similarly, Fang and Mao [17] consider BD-RIS with group/fully-connected architectures and derive a closed-form BD-RIS scattering matrix to maximize the sum of the users’ equivalent channel gains following the gradient decent approach and symmetric unitary projection, while regularized ZF precoding is considered at the BS. While relaxing the symmetry condition, the authors of [18] consider maximizing the sum rate and transforming the original problem into a multi-block optimization using auxiliary variables and solving it iteratively until convergence. In [19], Nerini et al. derive a closed-form BD-RIS scattering matrix that maximizes the channel gain of the single-user single-input-single-output (SISO) scenario and the solution is extended to the MU-MISO scenario. Furthermore, BD-RIS has been utilized in integrated sensing and communication (ISAC) systems [20, 21]. For instance, the authors of [20] jointly optimize the BS linear filter, and precoder, as well as the BD-RIS scattering matrix to maximize the sum rate while satisfying the sensing requirements. Chen and Mao [21] propose a two-stage design to jointly maximize the sum rate and minimize the largest eigenvalue of the Cramér-Rao bound matrix for multiple sensing targets.
@@ I-B Motivations and Contributions
Although a local optimal joint design was proposed in [18], the associated computational complexity of the proposed solution is significant. Therefore, a lower complexity two-stage design is proposed for multiuser systems in [17, 16]. Nonetheless, the focus in [17, 16, 19] is primarily on designing the BD-RIS scattering matrix to maximize the norm of the equivalent channel gains while the inter-user interference is neglected, and thus results in performance loss. Hence, it is more suitable for the point-to-point systems [22]. Therefore, this article focuses on efficiently and effectively addressing the interference management issue in multiuser systems by exploiting the interference management capability at fully-connected BD-RIS. The contributions of this work can be summarized as follows:
@@
    A two-stage design is considered for the BD-RIS scattering matrix and the BS precoding. Specifically, we analyze the received SINR at stage 1. We then design the BD-RIS scattering matrix to either maximize the sum received signal powers at the users following MRT structure, known as passive MRT, or to nullify the interference following ZF structure, known as passive interference nulling.
@@
    Closed-form BD-RIS scattering matrix expressions are presented for the passive MRT following the symmetric unitary projection [17], while an efficient AO algorithm is presented for the passive interference nulling. It is found that applying the symmetric unitary projection improves the received SINR at stage 1 for the passive MRT.
@@
    The computational complexity for both stages is analyzed considering the two BD-RIS designs and the BS processes. In addition, the channel state information requirements for each design are specified. For instance, it is found that the passive MRT eases the channel state information requirements which further simplifies the channel estimation.
@@ I-C Article Organization
The rest of the article is organized as follows. Sec. II introduces the system and channel models of the BD-RIS-enabled MU-MISO systems. Sec. III is the BD-RIS scattering matrix design for passive MRT and interference nulling. Sec. IV is the BS power allocation/precoding design. Sec. V is the numerical results and discussions. Finally, Sec. VI concludes the article with the main remarks and future work.
@@ II System And Channel Models
In a downlink MU-MISO system, the BS is equipped with $K$ transmit antennas to serve $K$ single antenna users through a fully-connected BD-RIS with $N$ REs. The precoded signal vector to be transmitted at the antenna ports of the BS can be expressed as $\mathbf{x}=\mathbf{Ps}$, where $\mathbf{s}\in\mathbb{C}^{K\times 1}$ is the information symbols vector with $\mathbb{E}\left[\mathbf{ss}^{H}\right]=\mathbf{I}$, and the precoding matrix is denoted by $\mathbf{P}\in\mathbb{C}^{K\times K}$ such that $\left\|\mathbf{P}\right\|^{2}_{F}=P_{\max}$ with $P_{\max}$ the maximum transmission power at the BS. Let the $k$th user be denoted as $U_{k}$, $k\in\{1,2,\ldots,K\}$, the received signal at $U_{k}$ is expressed as
@@
where $\mathbf{W}\in\mathbb{C}^{N\times K}$ is the BS-BD-RIS channel matrix, $\boldsymbol{\Theta}\in\mathbb{C}^{N\times N}$ is the BD-RIS scattering matrix, $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$ is BD-RIS-$U_{k}$ channel vector, and $n_{k}\sim\mathcal{CN}\left(0,N_{0}\right)$ is the additive white Gaussian noise (AWGN) with $N_{0}$ the power spectral density. In addition, $\boldsymbol{\Theta}$ belongs to the complex unitary symmetric set due to its lossless and reciprocal impedance network, i.e., $\boldsymbol{\Theta}\in\mathcal{S}_{\mathrm{loss}}=\{\boldsymbol{\Theta}|\boldsymbol{\Theta}\boldsymbol{\Theta}^{H}=\mathbf{I}\}$ and $\boldsymbol{\Theta}\in\mathcal{S}_{\mathrm{recip}}=\{\boldsymbol{\Theta}|\boldsymbol{\Theta}=\boldsymbol{\Theta}^{T}\}$ [13]. It is worth noting that the BS-$U_{k}$ channel is assumed to be blocked to highlight the performance gain of the BD-RIS [19]. To allow a systematic design of the scattering matrix, (1) can be written in a compact form such that the received signal vector is $\mathbf{r}=\mathbf{H}\boldsymbol{\Theta}\mathbf{Wx}+\mathbf{n}\in\mathbb{C}^{K\times 1}$, where $\mathbf{H}=\left[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}\right]^{T}\in\mathbb{C}^{K\times N}$ is the concatenated BD-RIS-$U_{k}$, $\forall k$ channel matrix, and $\mathbf{n}$ is the AWGN vector with independent and identically distributed (i.i.d) entries. Fig. 1 depicts the system block diagram.
@@
$\boldsymbol{\Theta}$ $=\boldsymbol{\Theta}^{T}$ (3c)
$\boldsymbol{\Theta}\boldsymbol{\Theta}^{H}$ $=\mathbf{I}$ (3d)
The objective function in ($P1$) shows a strong coupling between the precoding matrix and the scattering matrix, which makes it difficult to jointly optimize two matrices. Therefore, it is suggested to have a two-stage design. Unlike [17], our proposed design directly enhances the SINR by applying passive MRT/interference nulling at the BD-RIS. Such approach is motivated by the fact that utilizing ZF solely at the BS usually incurs power penalty because most of the power is wasted for channel inversion, resulting in having very small power gains [23]. Hence, due to the BD-RIS passive nature, enhancing the SINR by the BD-RIS is more effective than relying solely on the BS for precoding.

### v2

@@ Abstract
Beyond diagonal reconfigurable intelligent surfaces (BD-RIS) generalizes and goes beyond conventional diagonal reconfigurable intelligent surfaces (D-RIS) by interconnecting elements to generate beyond diagonal scattering matrices, which significantly strengthen the wireless channels. In this work, we use BD-RIS for passive multiuser beamforming in multiuser multiple-input-single-output (MU-MISO) systems. Specifically, we design the scattering matrix of BD-RIS to either maximize the sum received signal power at the users following maximum ratio transmission (MRT), or to nullify the interference at the users following zero forcing (ZF). To control the BD-RIS circuit topology complexity, we present the scattering matrix designs for the single/group/fully-connected BD-RIS architectures. Furthermore, we investigate uniform/optimized power allocation and ZF precoding at the base station (BS). Numerical results show that BD-RIS improves the interference nulling capability and sum rate with fewer reflecting elements (REs) compared to D-RIS. In addition, at moderate to high signal to noise ratios (SNRs), passive interference nulling reduces the complexity at the BS by relaxing the need for precoding or water-filling power allocation design. Furthermore, the passive MRT with ZF precoding achieves a tight sum rate performance to the joint design considering MU-MISO scenarios with many REs while maintaining low computational complexity and simplifying the channel estimation.
@@ I-A Related Work
RIS technologies have gained significant attention recently due to their potential to enhance the communication performance of wireless systems. Among these technologies, two prominent categories have been intensively investigated in the literature, which are diagonal reconfigurable intelligent surfaces (D-RIS) [6, 7, 8, 9, 10, 11, 12] and beyond diagonal reconfigurable intelligent surfaces (BD-RIS) [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. While these two categories are often discussed separately, it is important to note that BD-RIS can be viewed as a generalization of D-RIS. In other words, BD-RIS encompasses the characteristics of D-RIS as a special case, offering a broader and more flexible framework for reconfigurable intelligent surfaces. The key limitation in D-RIS is that it manipulates only the phase of the incident signal with $N$ reconfigurable impedances attached to $N$ REs making single-connected circuit topology. In contrast, BD-RIS manipulates both the phase and amplitude of the incident signal with a network of reconfigurable impedances that connect the $N$ REs together [13, 14]. By interconnecting all the REs to each other through $\frac{N(N+1)}{2}$ reconfigurable impedances, the fully-connected BD-RIS has been proposed, having the highest flexibility yet the highest circuit topology complexity. To balance the performance of the fully-connected BD-RIS with its circuit topology complexity, group-connected BD-RIS is introduced by Shen et al. [13]. In this architecture, the REs are divided into $G$ groups, each with a group size of $N_{G}=\frac{N}{G}$ REs. Within each group, the REs are interconnected using $\frac{N_{G}(N_{G}+1)}{2}$ reconfigurable impedances, reducing the total number of reconfigurable impedances to $\frac{N(N_{G}+1)}{2}$. Additionally, graph theory is utilized to design optimal architectures with lower circuit topology complexity such as forest-connected and tree-connected [15].
@@
Focusing on various RIS architectures, in existing literature, beamforming design for both D-RIS [6, 7, 8, 9, 10, 11, 12] and BD-RIS [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] aided multiuser multiple-input-single-output (MU-MISO) systems have been studied and explained in detail below.
@@
D-RIS: The authors of [6] minimize the transmit power at the base station (BS) while maintaining a certain signal to interference and noise ratio (SINR) level at the users by jointly optimizing the scattering matrix of D-RIS and the active precoder at the BS using the alternating optimization (AO) algorithm. Similarly, Yang et al. [7] achieve the same goal using passive maximum ratio transmission (MRT)/zero forcing (ZF) at the D-RIS while iteratively optimizing the power allocation at the BS. Considering the same setup, the authors of [8] use AO framework to maximize the minimum rate for a better user fairness. The authors of [9] derive a closed-form steering vectors-based multiuser linear precoder to null the known interference and minimize the unknown interference. Jiang et al. [10] use the AO framework to design the scattering matrix for interference nulling, achieving maximum degrees-of-freedom with high probability when the number of REs is $N>2K(K-1)$, with $K$ the number of users. The authors of [11] present a robust gradient-based meta-learning approach without pre-training to design the scattering matrix of the D-RIS and the precoder at the BS. Jangsher et al. [12] maximize the sum rate of a wireless unmanned aerial vehicle (UAV) with RIS considering energy efficiency constraints and practical limitations such as phase compensation and imperfect channel state information (CSI).
@@
BD-RIS: A joint design for the BD-RIS scattering matrix and BS precoder is considered in [16, 17, 18, 19, 20], unlike [21] which considers a two-stage design. The authors of [16, 17] maximize the sum rate of a multi-sector BD-RIS aided system by transforming the original problem into a multi-block optimization using auxiliary variables and solving it iteratively until convergence. In [18], a unified approach is proposed for various objectives such as energy efficiency maximization, energy consumption minimization and sum rate maximization, while satisfying all constraints on the BD-RIS scattering matrix with additional quality-of-service constraints. In [19], Nerini et al. derive a closed-form BD-RIS scattering matrix that maximizes the channel gain of the single-user single-input-single-output (SISO) scenario and the solution is extended to the MU-MISO scenario. Sun et al. [20] utilize the Takagi factorization-based symmetric unitary projection and AO to jointly design the BS precoder and the fully-connected BD-RIS scattering matrix. The BS precoder aims to minimize the transmitted power and satisfy certain SINR thresholds at the users, while the BD-RIS scattering matrix aims to maximize the equivalent channel’s squared trace magnitude. Fang and Mao [21] consider BD-RIS with group/fully-connected architectures and derive a closed-form BD-RIS scattering matrix to maximize the sum of the users’ equivalent channel gains following the gradient decent approach and the singular value decomposition (SVD)-based symmetric unitary projection, while regularized ZF precoding is considered at the BS.
Furthermore, BD-RIS has been investigated in cell-free massive multiple-input-multiple-output (MIMO) [22], and integrated sensing and communication (ISAC) systems [23, 24]. For instance, the authors of [22] employed BD-RIS in simultaneous wireless information and power transfer (SWIPT) systems to improve the wireless power transfer. The authors of [23] jointly optimize the BS linear filter, and precoder, as well as the BD-RIS scattering matrix to maximize the sum rate while satisfying the sensing requirements. Chen and Mao [24] propose a two-stage design to jointly maximize the sum rate and minimize the largest eigenvalue of the Cramér-Rao bound matrix for multiple sensing targets. While the aforementioned works consider reciprocal impedance networks, non-reciprocal single/group/multi-sector-connected BD-RIS architectures are proposed in [25, 26, 27] by relaxing the symmetry constraint.
@@ I-B Motivations and Contributions
Although a local optimal joint design was proposed in [16, 17], the associated computational complexity of the proposed solution is significant. Therefore, a lower complexity two-stage design is proposed for multiuser systems in [21]. Nonetheless, the focus in [21] is primarily on designing the BD-RIS scattering matrix to maximize the norm of the equivalent channel gains while the inter-user interference is neglected, and thus results in performance loss. Hence, it is more suitable for the point-to-point systems as shown in [28]. Therefore, this article focuses on efficiently and effectively addressing the interference management issue in multiuser systems by exploiting the interference management capability at the BD-RIS. The contributions of this work can be summarized as follows:
@@
    An efficient and low-complexity two-stage design is proposed for the BD-RIS scattering matrix and the BS precoding. Specifically, we analyze the received SINR at stage 1. We then design the BD-RIS scattering matrix to either maximize the sum received signal powers at the users following MRT structure, known as passive MRT, or to nullify the interference following ZF structure, known as passive interference nulling. The proposed BD-RIS scattering matrix design is general for the single/group/fully-connected BD-RIS architectures.
@@
    Closed-form BD-RIS scattering matrix expressions are presented for the passive MRT following the SVD-based symmetric unitary projection [21], while an efficient AO algorithm is presented for the passive interference nulling. It is found that applying the SVD-based symmetric unitary projection improves the received SINR at stage 1 for the passive MRT.
@@
    The computational complexity for both stages is analyzed considering the two BD-RIS scattering matrix designs, the three BD-RIS architectures, and the BS processes. In addition, the channel state information requirements for each design are specified. For instance, it is found that the passive MRT eases the channel state information requirements which further simplifies the channel estimation.
@@ I-C Article Organization
The rest of the article is organized as follows. Sec. II introduces the system and channel models of the BD-RIS-enabled MU-MISO systems. Sec. III is the BD-RIS scattering matrix design for passive MRT and interference nulling considering single/group/fully-connected BD-RIS architectures. Sec. IV is the BS power allocation/precoding design. Sec. V is the numerical results and discussions. Finally, Sec. VI concludes the article with the main remarks and future work. Interested readers are referred to the Matlab code for this article which is available on https://github.com/hmjasmi/bdris-mrt-null.
@@ II System and Channel Models
In a downlink MU-MISO system, the BS is equipped with $K$ transmit antennas to serve $K$ single antenna users through a BD-RIS with $N$ REs. The precoded signal vector to be transmitted at the antenna ports of the BS can be expressed as $\mathbf{x}=\mathbf{Ps}$, where $\mathbf{s}\in\mathbb{C}^{K\times 1}$ is the information symbols vector with $\mathbb{E}\left[\mathbf{ss}^{H}\right]=\mathbf{I}$, and the precoding matrix is denoted by $\mathbf{P}\in\mathbb{C}^{K\times K}$ such that $\left\|\mathbf{P}\right\|^{2}_{F}=P_{\max}$ with $P_{\max}$ the maximum transmission power at the BS. Let the $k$th user be denoted as $U_{k}$, $k\in\{1,2,\ldots,K\}$, the received signal at $U_{k}$ is expressed as
@@
where $\mathbf{W}\in\mathbb{C}^{N\times K}$ is the BS-BD-RIS channel matrix, $\boldsymbol{\Theta}\in\mathbb{C}^{N\times N}$ is the BD-RIS scattering matrix, $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$ is BD-RIS-$U_{k}$ channel vector, and $n_{k}\sim\mathcal{CN}\left(0,N_{0}\right)$ is the additive white Gaussian noise (AWGN) with $N_{0}$ the power spectral density. It is worth noting that the BS-$U_{k}$ channel is assumed to be blocked to highlight the performance gain of the BD-RIS [19]. To allow a systematic design of the scattering matrix, (1) can be written in a compact form such that the received signal vector is $\mathbf{r}=\mathbf{H}\boldsymbol{\Theta}\mathbf{Wx}+\mathbf{n}\in\mathbb{C}^{K\times 1}$, where $\mathbf{H}=\left[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}\right]^{T}\in\mathbb{C}^{K\times N}$ is the concatenated BD-RIS-$U_{k}$, $\forall k$ channel matrix, and $\mathbf{n}$ is the AWGN vector with independent and identically distributed (i.i.d) entries. Fig. 1 depicts the system block diagram.
@@
Depending on the circuit topology for the BD-RIS reconfigurable impedances network, three architectures are considered [13]:
  1. 1.
    Single-connected BD-RIS: The single-connected BD-RIS is equivalent to the conventional D-RIS. Each RE in this architecture is independently connected with a single grounded reconfigurable impedance without any interconnections among different REs. Consequently, the scattering matrix of the single-connected BD-RIS is a diagonal matrix, which gives rise to the term D-RIS representing the independent operation of each RE. The constraint on the scattering matrix of the single-connected BD-RIS is given as
    $\mathcal{S}_{\mathrm{SC}_{1}}=\left\{\boldsymbol{\Theta}:\left[\boldsymbol{\Theta}\right]_{i,j}=0,\,\forall i\neq j\right\}$ (2)
    where $i,j\in\{1,2,\ldots,N\}$. Considering purely reactive impedances, the scattering matrix of the lossless single-connected BD-RIS shall satisfy the unit-modulus constraint for all its elements, i.e.,
    $\mathcal{S}_{\mathrm{SC}_{2}}=\left\{\boldsymbol{\Theta}:\left|\left[\boldsymbol{\Theta}\right]_{i,j}\right|=1,\,\forall i=j\right\}$ (3)
  2. 2.
    Fully-connected BD-RIS: In the fully-connected BD-RIS architecture, each RE is interconnected with all other REs through a network of reconfigurable impedances, allowing waves impinging on one RE to be reflected by all other REs. Consequently, the scattering matrix of the fully-connected BD-RIS is a full, reciprocal matrix due to the linearity of the reconfigurable impedances network. Therefore, the symmetry constraint on the scattering matrix of the fully-connected BD-RIS is given as
    $\mathcal{S}_{\mathrm{FC}_{1}}=\left\{\boldsymbol{\Theta}:\boldsymbol{\Theta}=\boldsymbol{\Theta}^{T}\right\}$ (4)
    Considering purely reactive impedances, the scattering matrix of the lossless fully-connected BD-RIS shall satisfy the unitary constraint, i.e.,
    $\mathcal{S}_{\mathrm{FC}_{2}}=\left\{\boldsymbol{\Theta}:\boldsymbol{\Theta}\boldsymbol{\Theta}^{H}=\mathbf{I}\right\}$ (5)
  3. 3.
    Group-connected BD-RIS: The group-connected BD-RIS architecture divides the REs into $G$ groups, each with a group size of $N_{G}=\frac{N}{G}$ REs. Within each group, the REs are interconnected with all other intra-group REs through a network of reconfigurable impedances, allowing waves impinging on one RE to be reflected by all other intra-group REs only because inter-group REs are disconnected. Consequently, the scattering matrix of the group-connected BD-RIS is a block diagonal, with reciprocal block matrices due to the linearity of the reconfigurable impedances network in each group. Therefore, the symmetry constraint on each block of the scattering matrix of the group-connected BD-RIS is given as
    $\mathcal{S}_{\mathrm{GC}_{1}}=\bigl{\{}\boldsymbol{\Theta}:\boldsymbol{\Theta}=\mathrm{blkdiag}\left(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{G}\right),\,\\ \boldsymbol{\Theta}_{g}=\boldsymbol{\Theta}_{g}^{T},\,\forall g\bigr{\}}$ (6)
    where $\boldsymbol{\Theta}_{g}\in\mathbb{C}^{N_{G}\times N_{G}}$, $g\in\{1,2,\ldots,G\}$. Considering purely reactive impedances, the scattering matrix of the lossless group-connected BD-RIS shall satisfy the unitary constraint on each block $\boldsymbol{\Theta}_{g}$, i.e.,
    $\mathcal{S}_{\mathrm{GC}_{2}}=\bigl{\{}\boldsymbol{\Theta}:\boldsymbol{\Theta}_{g}\boldsymbol{\Theta}_{g}^{H}=\mathbf{I},\,\forall g\bigr{\}}$ (7)
    It is worth noting that when $G=1$ and $G=N$, the group-connected boils down to the fully-connected and single-connected architectures, respectively.
@@
$\boldsymbol{\Theta}$ $\in\mathcal{S}_{c_{1}}$ (9c)
$\boldsymbol{\Theta}$ $\in\mathcal{S}_{c_{2}}$ (9d)
where $c\in\{\mathrm{SC},\,\mathrm{FC},\,\mathrm{GC}\}$. The objective function shown in (9a) is non-convex and shows a strong coupling between the precoding matrix and the scattering matrix, which makes it difficult to jointly optimize the two matrices. The constraint shown in (9b) is non-convex. The constraint shown in (9c) stems from the BD-RIS circuit topology, which is non-convex in general. The constraint shown in (9d) is non-convex [17], and it stems from the lossless nature of the BD-RIS. It is widely known as the Stiefel manifold condition [18]. Therefore, it is suggested to have a two-stage design to decouple the two optimization variables. Unlike [21], our proposed design directly enhances the SINR by applying passive MRT/interference nulling at the BD-RIS. Such approach is motivated by the fact that utilizing ZF solely at the BS usually incurs power penalty because most of the power is wasted for channel inversion, resulting in having very small power gains [29]. Hence, due to the BD-RIS passive nature, enhancing the SINR by the BD-RIS is more effective than relying solely on the BS for precoding.